What mass of chromium metal is in a sample that contains 6.9 × 1025 formula units of ammonium dichromate, (NH4)2Cr2O7(s)?


1
Expert's answer
2021-11-11T01:34:03-0500

Moles of "(NH_4)_2Cr_2O_7=\frac{6.9\times 10^{25}}{6.022\times 10^{23}}=114.57 moles"

Moles of Cr "=2\times 114.57=229.15moles"

Mass of Cr"=229.16mol \times 51.99g/mol=1.2\times 10^4 g"